OpenStudy (anonymous):

Solve for x. State any restrictions on the variables. a(bx+2)=cx-12; A. x=(-12-2a)/(ab-c);ab≠c B. x=(-12+2a)/(c+ab);ab≠c C. x=(2a+12)/(c-ab);no restrictions D. x=(ab-c)/(-12-2a);a≠-1

OpenStudy (solomonzelman):

a(bx+2)=cx-12 abx+2a=cx-12 abx+2a-cx=-12 abx-cx=-12-2a x(ab-c)=-12-2a x= (-12-2a) / (ab-c)

OpenStudy (solomonzelman):

ab can't equal c

OpenStudy (anonymous):

thank you

OpenStudy (skullpatrol):

Did you learn anything @mikeiordan25 ?